OpenStudy (anonymous):

What are the first four terms of this multiplication pattern? The first term is 3, and then each term is multiplied by 5.

OpenStudy (anonymous):

Me will give a fan and a Medal ^-^

OpenStudy (anonymous):

3*5=15 15*5=75 75*5=375 375*5=1875 so 3, 15, 75,375,1875

OpenStudy (anonymous):

Thanks liliy

OpenStudy (anonymous):

you can also use the geometric series formula: an= a1* r ^(n-1)

OpenStudy (anonymous):

can you answer another one

OpenStudy (anonymous):

Glenn invested $18,000 at 5% interest compounded annually. How much interest will Glenn earn in 3 years?

OpenStudy (anonymous):

A. $2837.25 B. $20,837.25 C. $2340.23 D. $19,826.24

OpenStudy (anonymous):

i think its C or A

OpenStudy (anonymous):

use the compound interest formula:A=P*e^rt

OpenStudy (anonymous):

A=18,000*e^(.05*3)

OpenStudy (anonymous):

B.20,837.25

OpenStudy (anonymous):

P(1 + r/n)nt . this is the formula to plug it into: p is principle, r is rate, n is how many times per year. t is time